# Learning an Ensemble Token from Task-driven Priors in Facial Analysis

### v1

@@ 3.3 Decoder and inference
The $\hat{t}_{k}$ and $\hat{t}_{c}$ derived from the ET-Fuser are subjected to element-wise addition with the spatial embedding information $e_{i}$. The plenty feature embedding $\hat{e}_{c}$ is subsequently provided as input to the canonical decoder $d_{c}$. While a set of N decoders $D=\{d_{1},d_{2},...,d_{n-1},d_{c}\}$ is potentially learnable for other tasks, only the task-specific decoder $d_{c}$ undergoes fine-tuning to accommodate $\hat{e}_{c}$. This canonical, or selective, fine-tuning strategy restricts the trainable parameters to the ET-Fuser and the canonical decoder, thereby significantly reducing the computational resources required for both training and inference. Each decoder comprises a light multi-layer perceptron architecture tailored for classification and regression tasks. For the loss function, WingLoss [8] was employed for regression tasks, including age estimation and landmark detection, and cross-entropy loss was employed for classification tasks, including face emotion recognition and face recognition.
Figure 5: Inference phrase of decoupled encoder. The architectural parameters of all layers are maintained consistently with those employed during the training phase. Only the difference lies in the input provided to the ET-Fuser module, which is exclusively the canonical token.
During inference, only $t_{c}$ derived from $e_{c}$ are presented as input to the ET-Fuser as shown in Fig. 5. While the information density is lower than training, $t_{k}$ retains a sophisticated and entangled representation of the priors acquired during the training phase. The ET-Fuser’s training on ”thinned” information, facilitated by the drop mask, ensures that the inference process exhibits a high degree of fidelity to the training phrase. To facilitate the application of positional embeddings to $T_{\text{masked}}$, the task index corresponding to each encoder is preserved throughout the training phase. This preservation guarantees the consistent application of positional embeddings to corresponding token’s locations, even in scenarios where elements from $T\setminus{t_{c}}$ are absent in inference phrase. Canonical masking operations are omitted during inference as it identically equals to $\{\hat{t}_{k},\hat{t}_{c}\}$. In consequence, the exclusion of other $e_{i}$ results in a substantial reduction in computational overhead compared to ensemble learning, with the computational savings scaling linearly with the number of encoders.
@@ 4.1 Implementation details
The experimental evaluations were conducted on 2 NVIDIA Tesla H100 GPUs. For the backbone network architecture, MobileOne [40] was selected to assess real-time performance on resource-constrained platforms, such as embedded systems and hands-on devices, while ConvNeXt-b (CNN-based) [44] and Swin-b (ViT-based) [20] were employed to evaluate performance on desktop-level hardware. All input images were resized to 224$\times$224 pixels, with facial regions cropped using a YOLO-based [32] face detector. No facial alignment procedures were applied; however, task-specific pre-processing pipelines were implemented to ensure robust generalization by tasks. Specifically, for age estimation and face recognition tasks, the facial region was cropped to encompass both the facial area and hair, whereas for face landmark detection and face emotion recognition, a more tightly cropped region was adapted. A uniform batch size of 32 was maintained across all tasks. Given the head-tuning paradigm employed, which is analogous to linear probing, a learning rate of 1e-5 was adopted in trainable layers with warm-up cosine scheduler. In terms of computational cost, the ET-Fuser exhibited 21.02 MFLOPs during the inference phase, with a token dimension size of 1024. During the training phase, the computational cost increases proportionally with the number of tasks.
4.2 Datasets and prior tasks
@@
4.3 Pre-training the prior models
The experimental design was formulated to assess the variability in feature representations arising from disparities in model architecture and capacity. Accordingly, MobileOne-s0 (M), ConvNeXt V2-base (C), and Swin Transformer-base (S) were chosen as baselines, representing lightweight CNN, heavyweight CNN, and hierarchical ViT architectures, respectively. Each model underwent pre-training on four prior tasks, utilizing a simple MLP decoder head. In terms of computational cost, considering that MobileOne-s0 operates at about 275 MFLOPs, ET-Fuser’s 21.02 MFLOPs introduces a marginal computational burden. In comparison to other computationally demanding models, this overhead is effectively inconsequential.
For illustrative examples as experimental table, ”M (baseline)” denotes the architectural structure pre-trained on MobileOne, while ”C to M” denotes a configuration wherein ConvNeXt serves as the pre-trained model for the N-1 prior branches, and MobileOne is designated as the canonical branch for ensemble token learning. For visualizing the embeddings, we employed the t-SNE [38] to reduce the N-dimensional embeddings to two-dimensional embeddings in Fig. 6, Fig. 7, and Fig. 8.
4.4 Face landmark detection
  WFLW
Structure NME $\downarrow$ MAE $\downarrow$
M (baseline) 5.016 3.104
M to M 4.984 (-0.032) 3.084 (-0.020)
C to M 4.976 (-0.040) 3.080 (-0.024)
S to M 5.020 (+0.004) 3.102 (-0.002)
C (baseline) 4.398 2.735
M to C 4.360 (-0.038) 2.712 (-0.022)
C to C 4.350 (-0.048) 2.708 (-0.027)
S to C 4.360 (-0.038) 2.717 (-0.018)
S (baseline) 4.543 2.829
M to S 4.512 (-0.031) 2.804 (-0.025)
C to S 4.529 (-0.014) 2.811 (-0.018)
S to S 4.525 (-0.018) 2.806 (-0.023)
To assess face landmark detection performance on the WFLW dataset, normalized mean error (NME) and mean absolute error (MAE) on 98 landmark points were adopted as evaluation metrics. The NME was normalized by the inter-ocular distance, defined as the Euclidean distance between the lateral canthi of the left and right eyes. The MAE was scaled relative to the input size of 224. Although significant performance enhancements were not universally demonstrated in Tab. 1, all metrics exhibited a statistically significant improvement with the ET-Fuser methodology, with the exception of NME of ”S to M” with +0.004. Notably, the adoption of CNN-based task priors consistently yielded more pronounced performance compared to ViT-based priors.
4.5 Age estimation
(a) M (baseline)
(b) C to M (ours)
Figure 6: Visualization of embeddings (token) in UTK-face dataset. The color of each data point corresponds to the age represented in the right color-bar. Notably, data points that were distinct within the red ellipse in Fig. 6(a) are not observed in the ”C to M” data points in Fig. 6(b). This observation suggests that corner-case scenarios have been effectively assimilated through the ET-Fuser training phrase.
  UTK-face
Structure CS@5 $\uparrow$ MAE $\downarrow$
M (baseline) 66.79 4.568
M to M 67.29 (+0.50%) 4.548 (-0.020)
C to M 66.85 (+0.06%) 4.549 (-0.019)
S to M 67.24 (+0.45%) 4.554 (-0.014)
C (baseline) 68.32 4.316
M to C 68.66 (+0.34%) 4.285 (-0.031)
C to C 69.10 (+0.78%) 4.288 (-0.028)
S to C 69.30 (+0.98%) 4.277 (-0.039)
S (baseline) 67.49 4.412
M to S 67.53 (+0.04%) 4.402 (-0.010)
C to S 67.60 (+0.11%) 4.400 (-0.012)
S to S 67.49 ($\pm$0%) 4.402 (-0.010)
Table 2: Evaluation table on UTK-face dataset. The CS demonstrates performance improvement with increasing values, denoted by an up-arrow. Conversely, the MAE demonstrates performance improvement with decreasing values, denoted by a down-arrow.
For the evaluation of age estimation performance, MAE and cumulative score (CS) metrics were employed on the UTK-face dataset. The MAE was computed as the absolute deviation between predicted and ground truth ages, spanning the range of 1 to 116 years. The CS metric was determined as the percentage of predictions falling within $\pm$2 years, a tolerance level of 5. Consistent with observations in landmark detection, MAE exhibited marginal performance enhancements, lacking substantial improvements in Tab. 4. However, the CS metric demonstrated statistically significant performance gains. Notably, the adaptation of task priors among high-capacity models within the ConvNeXt architecture and Swin Transformer yielded superior performance improvements, ”C to C” with +0.78% and ”S to C” +0.98%, compared to alternative architectures, which achieves substantial performance gains relative to ET-Fuser’s minimal computational overhead.
4.6 Face emotion recognition
(a) M (baseline)
(b) C to M (ours)
Figure 7:  Visualization of embeddings in Affect-Net-HQ dataset. Each color delineates a unique category within the facial emotion classification. Quantitative analysis of intra-class variance, conducted within a the consistent spatial range, revealed a variance of 0.2641 on Fig. 7(a), while the Fig. 7(b) demonstrated a reduced variance of 0.2406, corresponding to an approximate 9% squeeze in intra-class dispersion.
  Affect-Net-HQ
Structure Accuracy $\uparrow$ F1-score $\uparrow$
M (baseline) 0.7261 0.7023
M to M 0.7309 (+0.48%) 0.7052 (+0.29%)
C to M 0.7299 (+0.38%) 0.7040 (+0.17%)
S to M 0.7323 (+0.62%) 0.7063 (+0.40%)
C (baseline) 0.7467 0.7233
M to C 0.7434 (-0.33%) 0.7204 (-0.29%)
C to C 0.7465 (-0.02%) 0.7235 (+0.02%)
S to C 0.7472 (+0.05%) 0.7245 (+0.12%)
S (baseline) 0.7390 0.7176
M to S 0.7513 (+1.23%) 0.7276 (+1.00%)
C to S 0.7520 (+1.30%) 0.7277 (+1.01%)
S to S 0.7501 (+1.11%) 0.7263 (+0.87%)
Table 3: Evaluation table on Affect-Net-HQ dataset.
For the classification task of face emotion recognition, we employed standard evaluation metrics, accuracy and F1-score. As shown in Table 3, all methodologies, with the exception of ConvNeXt, exhibited performance enhancements. We posit that this phenomenon is attributable to the inherent high capacity of ”C” for the emotion recognition task relative to the other task priors. Notably, despite ”S” demonstrating significantly lower accuracy compared to ”C” (0.7390 and 0.7176 vs. 0.7467 and 0.7233), the application of ET-Fuser in the ”C to S” outperformed with superior performance (0.7520 and 0.7277). This outcome is particularly compelling, as the observed performance gains in the ”S to C”, where the task priors are reversed to ”C to S”, suggest a synergistic relationship between ConvNeXt and Swin for face emotion recognition.
4.7 Face recognition
@@
Figure 8: Visualization of embeddings in CFP-dataset. Each color delineates a unique category within individual person identity. The elliptical contours visually depict the intra-class variance for each respective class. As corroborated by Tab. 4, the intra-class variance and the distinction of classes is significantly improved in the Fig. 8(b) compared to the Fig. 8(a)
@@
For face recognition evaluation, we employed accuracy, F1-score, intra-class variance, and true acceptance rate (TAR), a metric commonly employed in biometric recognition. Validation on the CFP dataset, which contains a high proportion of challenging examples, was performed at a false acceptance rate (FAR) of 0.1%. Identity classification demonstrated performance gains in both accuracy and F1-score within CNN architectures. The Swin architecture exhibited a mixed pattern of performance gains and losses, which we attribute to the weak inductive biases associated with ViT when applied to recognition tasks. Notably, MobileOne with ET-Fuser exhibited dramatic performance improvements, with intra-class variance reduced by over half and TAR gains over 10%. While ”S to C” and ”S to S” exhibited minor decreases in intra-class variance, TAR exhibited minor increases. This observation suggests improved discriminative power and feature representation within the inter-class space.
5 Limitations and future works

### v2

@@ 3.4 Decoder and Inference
The $\hat{t}_{k}$ and $\hat{t}_{c}$ derived from the KT-Adapter are subjected to element-wise addition with the spatial embedding information $e_{i}$. The feature embedding $\hat{e}_{c}$ is subsequently provided as input to the canonical decoder $d_{c}$. While a set of N decoders $D=\{d_{1},d_{2},...,d_{n-1},d_{c}\}$ is potentially learnable for other tasks, only the task-specific decoder $d_{c}$ undergoes fine-tuning to accommodate $\hat{e}_{c}$. This canonical, or selective, fine-tuning strategy restricts the trainable parameters to the KT-Adapter and the canonical decoder, thereby significantly reducing the computational resources required for both training and inference. Each decoder comprises a light multi-layer perceptron architecture tailored for classification and regression tasks. For the loss function, WingLoss [6] was employed for regression tasks, including age estimation and landmark detection, and cross-entropy loss was employed for classification tasks, including face emotion recognition and face recognition.
Figure 5: Inference phase of decoupled encoder. The architectural parameters of all layers are maintained consistently with those employed during the training phase. Only the difference lies in the input provided to the KT-Adapter module, which is exclusively the canonical token.
During inference, only $t_{c}$ derived from $e_{c}$ are presented as input to the KT-Adapter as shown in  5. While the information density is lower than training, $t_{k}$ retains a sophisticated and entangled representation of the priors acquired during the training phase. The KT-Adapter’s training on ”thinned” information, facilitated by the drop mask, ensures that the inference process exhibits a high degree of fidelity to the training phase. To facilitate the application of positional embeddings to $T_{\text{masked}}$, the task index corresponding to each encoder is preserved throughout the training phase. This preservation guarantees the consistent forward pass instead of positional embeddings to corresponding token’s locations, even in scenarios where elements from $T\setminus{t_{c}}$ are absent in inference phase. Canonical masking operations are omitted during inference as it identically equals to $\{\hat{t}_{k},\hat{t}_{c}\}$. In consequence, the exclusion of other $e_{i}$ results in a substantial reduction in computational overhead compared to knowledge learning, with the computational savings scaling linearly with the number of encoders.
@@ 4.1 Implementation Details
The experimental evaluations were conducted on 2 NVIDIA Tesla H100 GPUs. For the backbone network architecture, MobileOne [37] was selected to assess real-time performance on resource-constrained platforms, such as embedded systems and handheld devices, while ConvNeXt-b (CNN-based) [41] and Swin-b (ViT-based) [18] were employed to evaluate performance on desktop-level hardware. All input images were resized to $\times224$ pixels, with facial regions cropped using a YOLO-based [30] face detector. No facial alignment procedures were applied; however, task-specific pre-processing pipelines were implemented to ensure robust generalization by tasks. Specifically, for age estimation and face recognition tasks, the facial region was cropped to encompass both the facial area and hair, whereas for face landmark detection and face emotion recognition, a more tightly cropped region was adapted. A uniform batch size of 32 was maintained across all tasks. Given the head-tuning paradigm employed, which is analogous to linear probing, a learning rate of 1e-5 was adopted in trainable layers with warm-up cosine scheduler. In terms of computational cost, the KT-Adapter exhibited 21.02 MFLOPs during the inference phase, with a token dimension size of 1024. During the training phase, the computational cost increases proportionally with the number of tasks.
4.2 Datasets
@@
Figure 6: Visualization of embeddings in CFP-dataset. Each color delineates a unique category within individual person identity. The elliptical contours visually depict the intra-class variance for each respective class. As corroborated by  1, the intra-class variance and the distinction of classes is significantly improved in the  6(b) compared to the  6(a)
4.3 Pre-training the Prior Models
The experimental design was formulated to assess the variability in feature representations arising from disparities in model architecture and capacity. Accordingly, MobileOne-s0 (M), ConvNeXt V2-base (C), and Swin Transformer-base (S) were chosen as baselines, representing lightweight CNN, heavyweight CNN, and hierarchical ViT architectures, respectively. Each model underwent pre-training on four prior tasks, utilizing a simple MLP decoder head. In terms of computational cost, considering that MobileOne-s0 operates at about 275 MFLOPs, KT-Adapter’s 21.02 MFLOPs introduces a marginal computational burden. In comparison to other computationally demanding models, this overhead is effectively inconsequential.
For illustrative examples as experimental table, ”M (baseline)” denotes the architectural structure pre-trained on MobileOne, while ”C to M” denotes a configuration wherein ConvNeXt serves as the pre-trained model for the N-1 prior branches, and MobileOne is designated as the canonical branch for knowledge token learning. For visualizing the embeddings, we employed the t-SNE [35] to reduce the N-dimensional embeddings to two-dimensional embeddings in  7 and  6.
4.4 Evaluation per Datasets
@@
For face recognition evaluation, we employed accuracy, F1-score, intra-class variance, and true acceptance rate (TAR), a metric commonly employed in biometric recognition. Validation on the CFP dataset, which contains a high proportion of challenging examples, was performed at a false acceptance rate (FAR) of 0.1%. Identity classification demonstrated performance gains in both accuracy and F1-score within CNN architectures. The Swin architecture exhibited a mixed pattern of performance gains and losses, which we attribute to the weak inductive biases associated with ViT when applied to recognition tasks. Notably, MobileOne with KT-Adapter exhibited dramatic performance improvements, with intra-class variance reduced by over half and TAR gains over 10%. While ”S to C” and ”S to S” exhibited minor decreases in intra-class variance, TAR exhibited minor increases. This observation suggests improved discriminative power and feature representation within the inter-class space.
  WFLW
Structure NME $\downarrow$ MAE $\downarrow$
M (baseline) 5.016 3.104
M to M 4.984 (-0.032) 3.084 (-0.020)
C to M 4.976 (-0.040) 3.080 (-0.024)
S to M 5.020 (+0.004) 3.102 (-0.002)
C (baseline) 4.398 2.735
M to C 4.360 (-0.038) 2.712 (-0.022)
C to C 4.350 (-0.048) 2.708 (-0.027)
S to C 4.360 (-0.038) 2.717 (-0.018)
S (baseline) 4.543 2.829
M to S 4.512 (-0.031) 2.804 (-0.025)
C to S 4.529 (-0.014) 2.811 (-0.018)
S to S 4.525 (-0.018) 2.806 (-0.023)
To assess face landmark detection performance on the WFLW dataset, normalized mean error (NME) and mean absolute error (MAE) on 98 landmark points were adopted as evaluation metrics. The NME was normalized by the inter-ocular distance, defined as the Euclidean distance between the lateral canthi of the left and right eyes. The MAE was scaled relative to the input size of 224. Although significant performance enhancements were not universally demonstrated in  2, all metrics exhibited a statistically significant improvement with the KT-Adapter methodology, with the exception of NME of ”S to M” with +0.004. Notably, the adoption of CNN-based task priors consistently yielded more pronounced performance compared to ViT-based priors.
(a) M (baseline)
(b) C to M (ours)
Figure 7: Visualization of embeddings (token) in UTK-face dataset. The color of each data point corresponds to the age represented in the right color-bar. Notably, data points that were distinct within the red ellipse in  7(a) are not observed in the ”C to M” data points in  7(b). This observation suggests that the KT-Adapter not only assimilate the corner-case scenarios but also represent the feature clustering with adaptable transforms.
  UTK-face
Structure CS@5 $\uparrow$ MAE $\downarrow$
M (baseline) 66.79 4.568
M to M 67.29 (+0.50%) 4.548 (-0.020)
C to M 66.85 (+0.06%) 4.549 (-0.019)
S to M 67.24 (+0.45%) 4.554 (-0.014)
C (baseline) 68.32 4.316
M to C 68.66 (+0.34%) 4.285 (-0.031)
C to C 69.10 (+0.78%) 4.288 (-0.028)
S to C 69.30 (+0.98%) 4.277 (-0.039)
S (baseline) 67.49 4.412
M to S 67.53 (+0.04%) 4.402 (-0.010)
C to S 67.60 (+0.11%) 4.400 (-0.012)
S to S 67.49 ($\pm$0%) 4.402 (-0.010)
Table 3: Evaluation table on UTK-face dataset.
For the evaluation of age estimation performance, MAE and cumulative score (CS) metrics were employed on the UTK-face dataset. The MAE was computed as the absolute deviation between predicted and ground truth ages, spanning the range of 1 to 116 years. The CS metric was determined as the percentage of predictions falling within $\pm$2 years, a tolerance level of 5. Consistent with observations in landmark detection, MAE exhibited marginal performance enhancements, lacking substantial improvements in  1. However, the CS metric demonstrated statistically significant performance gains. Notably, the adaptation of task priors among high-capacity models within the ConvNeXt architecture and Swin Transformer yielded superior performance improvements, ”C to C” with +0.78% and ”S to C” +0.98%, compared to alternative architectures, which achieves substantial performance gains relative to KT-Adapter’s minimal computational overhead.
  Affect-Net-HQ
Structure Accuracy $\uparrow$ F1-score $\uparrow$
M (baseline) 0.7261 0.7023
M to M 0.7309 (+0.48%) 0.7052 (+0.29%)
C to M 0.7299 (+0.38%) 0.7040 (+0.17%)
S to M 0.7323 (+0.62%) 0.7063 (+0.40%)
C (baseline) 0.7467 0.7233
M to C 0.7434 (-0.33%) 0.7204 (-0.29%)
C to C 0.7465 (-0.02%) 0.7235 (+0.02%)
S to C 0.7472 (+0.05%) 0.7245 (+0.12%)
S (baseline) 0.7390 0.7176
M to S 0.7513 (+1.23%) 0.7276 (+1.00%)
C to S 0.7520 (+1.30%) 0.7277 (+1.01%)
S to S 0.7501 (+1.11%) 0.7263 (+0.87%)
Table 4: Evaluation table on Affect-Net-HQ dataset.
For the classification task of face emotion recognition, we employed standard evaluation metrics, accuracy and F1-score. As shown in 4, all methodologies, with the exception of ConvNeXt, exhibited performance enhancements. We posit that this phenomenon is attributable to the inherent high capacity of ”C” for the emotion recognition task relative to the other task priors. Notably, despite ”S” demonstrating significantly lower accuracy compared to ”C” (0.7390 and 0.7176 vs. 0.7467 and 0.7233), the application of KT-Adapter in the ”C to S” outperformed with superior performance (0.7520 and 0.7277). This outcome is particularly compelling, as the observed performance gains in the ”S to C”, where the task priors are reversed to ”C to S”, suggest a synergistic relationship between ConvNeXt and Swin for face emotion recognition.
  WFLW UTK-face Affect-Net-HQ CFP-dataset
Structure NME $\downarrow$ MAE $\downarrow$ CS@5 $\uparrow$ MAE $\downarrow$ Accuracy $\uparrow$ F1-score $\uparrow$ intra-C-var $\downarrow$ TAR@FAR=0.1% $\uparrow$
M (baseline) 5.016 3.104 66.79 4.568 72.61 70.23 0.6310 36.78
1 encoder 4.985 (-0.030) 3.084 (-0.020) 66.28 (+0.49%) 4.548 (-0.019) 73.08 (+0.47%) 70.50 (+0.28%) 0.4627 (-0.168) 49.44 (+12.67%)
2 encoders 4.976 (-0.039) 3.080 (-0.024) 66.85 (+0.05%) 4.550 (-0.017) 72.98 (+0.37%) 70.39 (+0.17%) 0.6090 (-0.021) 49.13 (+12.36%)
3 encoders(proposed) 4.976 (-0.040) 3.080 (-0.024) 67.29 (+0.50%) 4.548 (-0.020) 73.09 (+0.48%) 70.63 (+0.40%) 0.3186 (-0.312) 52.48 (+15.70%)
C (baseline) 4.398 2.735 68.32 4.316 74.67 72.33 0.1714 76.45
1 encoder 4.361 (-0.036) 2.713 (-0.021) 68.65 (+0.32%) 4.285 (-0.030) 74.34 (-0.33%) 72.04 (-0.29%) 0.1633 (-0.008) 78.91 (+2.47%)
2 encoders 4.352 (-0.045) 2.709 (-0.025) 69.05 (+0.72%) 4.289 (-0.026) 74.65 (-0.02%) 72.34 (+0.02%) 0.1672 (-0.004) 79.14 (+2.69%)
3 encoders(proposed) 4.350 (-0.048) 2.708 (-0.027) 69.30 (+0.98%) 4.277 (-0.039) 74.72 (+0.05%) 72.45 (+0.12%) 0.1632 (-0.008) 79.34 (+2.89%)
S (baseline) 4.543 2.829 67.49 4.412 73.90 71.76 0.4580 64.88
1 encoder 4.513 (-0.029) 2.804 (-0.024) 67.48 (-0.10%) 4.402 (-0.010) 75.09 (+1.19%) 72.74 (+0.99%) 0.3497 (-0.108) 69.11 (+4.23%)
2 encoders 4.530 (-0.012) 2.812 (-0.016) 67.60 (+0.10%) 4.401 (-0.011) 75.15 (+1.26%) 72.71 (+0.96%) 0.4523 (-0.005) 63.12 (-1.75%)
3 encoders(proposed) 4.512 (-0.031) 2.804 (-0.025) 67.60 (+0.11%) 4.400 (-0.012) 75.20 (+1.30%) 72.77 (+1.01%) 0.3493 (-0.108) 69.21 (+4.33%)
Table 5: Ablation study on parameter N, the number of encoders. We reported the results using the best-performing encoder among those sharing the same architecture type and encoder count. For example, in the face recognition task, the best performance was achieved when using encoders on age estimation and face landmark detection. We observed that increasing the number of prior encoders generally led to a linear improvement in performance.
       Structure Affect-Net-HQ UTK-face
DM CM PE Acc $\uparrow$ F1 $\uparrow$ CS@5 $\uparrow$ MAE $\downarrow$
0.7022 0.6677 62.50 4.771
0.7090 0.6998 66.00 4.601
0.7095 0.7002 65.89 4.588
0.7152 0.7020 66.51 4.599
0.7199 0.7022 66.51 4.598
0.7283 0.7041 66.87 4.550
0.7299 0.7040 66.85 4.549
Table 6: Ablation study on proposed modules.
4.5 Ablation Studies
An ablation study was conducted to assess the effects of the number of encoders N in  5, the use of drop mask ”DM”, canonical mask ”CM”, and the positional encoding ”PE” in  6. Evaluation metrics included accuracy and F1-score for facial emotion recognition, and CS@5 and MAE for age estimation. The use of the drop mask alone yielded improvements compared to the baseline model. Applying the canonical mask alone also yielded performance gains. Notably, a synergistic effect was observed when both the drop mask and canonical mask were employed simultaneously. As discussed in Section  3.4, this performance improvement indicates that structural misalignment between the training and inference phases plays a significant role in KT-Adapter. Interestingly, incorporating positional encoding into the prior tokens had only a minimal impact on performance. Considering the computational efficiency benefits of drop mask and canonical mask in training, we exclude positional encoding to reduce computational overhead.
5 Conclusion